40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 18 12:27:42 2019, gautam, Update, General, IFO status IMG_7343.JPGITMXOL.png
    Reply  Mon Mar 18 17:06:12 2019, gautam, Update, Optical Levers, ITMY HeNe replaced OLRIN.pdfOL_PIT.pdfOL_YAW.pdf
    Reply  Tue May 28 17:41:13 2019, gautam, Update, General, IFO status IMG_7444.JPG
Message ID: 14483     Entry time: Mon Mar 18 12:27:42 2019     Reply to this: 14484   14642
Author: gautam 
Type: Update 
Category: General 
Subject: IFO status 
  1. c1iscaux2 VME crate is damaged - see Attachment #1. 
    • It is not generating the 12V supply voltage, and so nothing in the crate works.
    • Tried resetting via front panel button, power cycling by removing power cable on rear, all to no effect.
    • Tried pulling out all cards and checking if there was an internal short that was causing the failure - looks like the problem is with the crate itself.
    • Not sure how long this machine has been unresponsive as we don't have any readback of the status of the eurocrate machines.
    • Not a showstopper, mainly we can't control the whitening settings for AS55, REFL55, REFL165 and ALSY. 
    • Acromag installation schedule should be accelerated.
    • * Koji reminded me that \text{VME crate} \ \neq \ \text{eurocrate}. The former is what is used for the slow machines, the latter is what is used for holding the iLIGO style electronics boards.
  2. ITMX oplev is dead - see Attachment #2.
    • Lasted ~3 years (installed March 2016).
    • I confirmed that no light is coming out of the laser head on the optical table.
    • I'll ask Chub to replace it this afternoon.
  3. c1susaux is unresponsive
    • I didn't reboot it as I didn't want to spend some hours freeing ITMY. 
    • At some point we will have to bite the bullet and do it.
  4. Input pointing is still not stable
    • I aligned the input pointing using TT1/TT2 to maximize TRX/TRY before lunch, but in 1 hour, the pointing has already drifted.
  5. POX/POY locking is working okay. TRX has large low-frequency fluctuations because of ITMX not having an Oplev servo, should be rectified once we swap out the HeNe.

The goal for this week is to test out the ALS system, so this is kind of a workable state since POX/POY locking is working. But the number of broken things is accumulating fast.

Attachment 1: IMG_7343.JPG  1.767 MB  Uploaded Mon Mar 18 13:32:25 2019  | Hide | Hide all
IMG_7343.JPG
Attachment 2: ITMXOL.png  27 kB  Uploaded Mon Mar 18 13:34:30 2019  | Hide | Hide all
ITMXOL.png
ELOG V3.1.3-